OpenStudy (anonymous):

can someone please help me with these equations of definite integrals in integration? ∫ x/sqrt 2x - 1 dx from 5 to 1 ∫ e^3/x//x^2 dx from 3 to 1 -- this is e to the power 3 over x, all over x squared. I hope that makes sense.

OpenStudy (anonymous):

\[\int\limits_{3}^{1} (e^3/x)/x^2 \] this what you mean?

OpenStudy (anonymous):

kinda of... but its e to the power 3/x.

OpenStudy (anonymous):

ln(2)/4 - ln(18)/4 - 2/9

OpenStudy (anonymous):

\[\int\limits_{5}^{1}x/(2x-1)^2\] =\[\ln(2)/4 - \ln(18)/4 - 2/9 \]

OpenStudy (anonymous):

regarding e^(3/x)/x^2 the derivative of 3/x is is -3/x^2 so let use substitution u=3/x du= -3/x^2 now we can do integral e^(u)*du take out the -1/3 ofcouse. so you should have \[-1/3\int\limits_{?}^{?} e^udu \] that should push you thou now sure about my answer:D but hope i helped

OpenStudy (anonymous):

thank you... any help is appreciated!! :)

OpenStudy (anonymous):

lol